# ModernVBERT: Towards Smaller Visual Document Retrievers

### v1

@@ A.3.1 Modality Alignment Mixture
For our modality alignment trainings, we rely on The Cauldron dataset (Laurençon et al., 2024b) and its Docmatix extension (Laurençon et al., 2024a). Table 5 provides further details on the constitution of this dataset.
@@
Table 5: Aggregated statistics of modality alignment datasets from The Cauldron 2 (Laurençon et al., 2024c) and Docmatix (Laurençon et al., 2024a), showing image counts, QA pairs, token counts, and the proportional contribution of each subsection to the overall mixture.
@@ A.3.2 NatCap
To enrich our contrastive learning data mixture, we construct NatCap (Natural Captions), a large-scale dataset containing around 333000 contextualized image–caption pairs. This dataset is created by generating synthetic captions, along with cross-class and in-class discriminative tags, from existing image classification datasets (see Table 6). For this purpose, we leverage Gemini-flash-2.5101010https://ai.google.dev/gemini-api/docs/models?hl=fr#gemini-2.5-flash which produces captions conditioned on both the image content and the accompanying dataset metadata, as illustrated in Figure 6. We detail the prompt below.
@@ A.3.3 Contrastive Training Mix
ColPali (Faysse et al., 2025) Query–Document images for visual retrieval 118k 1
MSCOCO (Lin et al., 2014) Natural images with human-written captions 118k 1
@@
RLHN (Thakur et al., 2025) Text–text pairs for complex retrieval 680k 1
@@
ColPali (Faysse et al., 2025) Query–Document images for visual retrieval 118k 3
RLHN (Thakur et al., 2025) Text–text pairs for complex retrieval 300k 3
@@ Appendix B Baselines Details
MoCa-3B (Chen et al., 2025). A modality-aware continual pretraining model that transforms a causal vision-language model into a bidirectional multimodal embedding model, using interleaved image-text reconstruction and contrastive alignment to support cross-modal retrieval.
GME-Qwen2 (Zhang et al., 2025a). A unified multimodal embedder built on Qwen2-VL (Wang et al., 2024), which produces shared embedding representations across text, image, and fused input modalities, enabling universal multimodal retrieval.
VLM2Vec (Jiang et al., 2025). A method that trains a vision-language encoder by converting a VLM through extensive contrastive post-training. Flagship model is based on the model Phi-3.5 (Abdin et al., 2024).
E5-V (Jiang et al., 2024). An adaptation of the E5 embedding approach to multimodal models: it trains only on text pairs yet bridges the modality gap to handle image inputs, reducing cost while achieving universal embeddings.
ColPali (Faysse et al., 2025). A vision-based document retrieval model that processes document pages as images (no OCR) and produces multi-vector embeddings via a late-interaction mechanism over PaliGemma (Beyer et al., 2024), enabling efficient and accurate retrieval.
ColQwen2.5 (Faysse et al., 2025). An extension of ColPali (Faysse et al., 2025) using Qwen2-VL (Wang et al., 2024) as the backbone, carrying forward the late interaction retrieval paradigm over page image embeddings, capturing layout and textual context without OCR.
Jina-v4 (Günther et al., 2025). A multimodal embedding model combining visual and textual inputs with support for multi-vector (late interaction) embeddings, using adapters over a unified backbone to excel on visually rich document retrieval.
NemoRetriever (Xu et al., 2025). An LI retriever that combines vision-language embeddings with a ColEmbed design, enabling high performance on visual document retrieval with structured patch matching and efficient similarity.
Jina CLIP (Koukounas et al., 2024). A smaller scale vision-language model using CLIP embeddings, applied to document retrieval tasks; although not LI, it offers a lightweight multimodal baseline.
BGE Visualized M3 (Zhou et al., 2024). A vision-enhanced version of BGE M3 (Chen et al., 2024) that supports visual inputs and extends embedding models into multimodal domains.
SigLIP2-L-512/16 (Tschannen et al., 2025). A multilingual vision-language bi-encoder model, which combines image and text modalities to yield unified embeddings across languages. This configuration handles images of 512x512 pixels and create subpatches of 16x16 pixels.
ColFlor (Masry & Hoque, 2024). A lightweight OCR-free visual document retriever with only 174M parameters built over Florence-2 and DaViT, delivering strong performance near ColPali with much lower computational cost and much faster encoding.
@@ Appendix C Additional Ablations
C.1 Scaling Dynamics of Attention Masks
@@
C.2 Bridging the Gap with Longer Contrastive Training
@@
C.2.1 Optimal Text-To-Image Ratio for Document Retrieval
@@
C.3 Late Interaction for Non-Documental Retrieval
@@
We want to study if the multi-vector gains transfer to non-documental retrieval. To do so, we contrastively post-train our base model on our generalist post-training mix presented in Table 7. The late interaction generalist exhibits superior performance in retrieval setting, improving its single-vector performance by +20.2% (11.5 points), matching the performance of substantially larger VLM-based retrievers like E5-V (8.3B parameters, 67.5 points) and surpassing dual encoders like SigLIP (882M parameters, 56.7 points). This matches the capabilities observed in Section 3.1 for documental settings for models with native bidirectional attention, extending it to natural image tasks. This result extends the prevailing understanding from the document retrieval community, where the superiority of late-interaction is well-documented (Khattab & Zaharia (2020), Chaffin (2025), Faysse et al. (2025)). While this performance gap is widely accepted for document retrieval, its applicability to caption matching tasks has not really been addressed. Our findings provide strong evidence that the fine-grained matching capabilities of late-interaction models are a key driver of performance in this domain too.
C.3.1 Model Merging
Our contrastive learning stage provides direct performance trade-offs on different tasks. Following recent trends, we evaluate how model merging techniques allow to mitigate performance degradation on specific tasks, while maintaining the performance enabled by the contrastive training (Sung et al., 2023; Dziadzio et al., 2024; Li et al., 2024; Zhang et al., 2025b). We merge our ablation model after modality alignment with the checkpoint after the full contrastive learning with two methods: SLERP (Ilharco et al., 2022) and average merging (Shoemake, 1985). For SLERP, we compare three values for the $\lambda$ coefficient ($0.25$, $0.5$, $0.75$). Figure 10 displays the the trends with the best method (SLERP, $\lambda=0.75$). As we can see, the merged model mitigates the performance drop in Image/Caption Retrieval tasks, while maintaining significant gains on Image Classification tasks. However, merging strongly degrades performance on Document Retrieval, showing that benefits of merging embedding models are task-dependent.
@@ C.3.1 Model Merging
C.3.2 Curriculum For Document Retriever Contrastive Post-Training
@@
C.4 Text-Only Retrieval

### v2

@@ A.3.1 Modality Alignment Mixture
For our modality alignment trainings, we rely on The Cauldron dataset (laurencon_what_2024) and its Docmatix extension (laurençon2024building). Table 5 provides further details on the constitution of this dataset.
@@
Table 5: Aggregated statistics of modality alignment datasets from The Cauldron 2 (laurençon2024mattersbuildingvisionlanguagemodels) and Docmatix (laurençon2024building), showing image counts, QA pairs, token counts, and the proportional contribution of each subsection to the overall mixture.
@@ A.3.2 NatCap
To enrich our contrastive learning data mixture, we construct NatCap (Natural Captions), a large-scale dataset containing around 333000 contextualized image–caption pairs. This dataset is created by generating synthetic captions, along with cross-class and in-class discriminative tags, from existing image classification datasets (see Table 6). For this purpose, we leverage Gemini-flash-2.5131313https://ai.google.dev/gemini-api/docs/models?hl=fr#gemini-2.5-flash which produces captions conditioned on both the image content and the accompanying dataset metadata, as illustrated in Figure 6. We detail the prompt below.
@@ A.3.3 Contrastive Training Mix
ColPali (faysse2025colpaliefficientdocumentretrieval) Query–Document images for visual retrieval 118k 1
MSCOCO (lin_microsoft_2014) Natural images with human-written captions 118k 1
@@
RLHN (thakur2025rlhn) Text–text pairs for complex retrieval 680k 1
@@
ColPali (faysse2025colpaliefficientdocumentretrieval) Query–Document images for visual retrieval 118k 3
RLHN (thakur2025rlhn) Text–text pairs for complex retrieval 300k 3
@@ Appendix B Baselines Details
MoCa-3B (chen2025mocamodalityawarecontinualpretraining). A modality-aware continual pretraining model that transforms a causal vision-language model into a bidirectional multimodal embedding model, using interleaved image-text reconstruction and contrastive alignment to support cross-modal retrieval.
GME-Qwen2 (zhang2025gmeimprovinguniversalmultimodal). A unified multimodal embedder built on Qwen2-VL (wang2024qwen2vlenhancingvisionlanguagemodels), which produces shared embedding representations across text, image, and fused input modalities, enabling universal multimodal retrieval.
VLM2Vec (jiang2025vlm2vectrainingvisionlanguagemodels). A method that trains a vision-language encoder by converting a VLM through extensive contrastive post-training. Flagship model is based on the model Phi-3.5 (abdin_phi-3_2024).
E5-V (jiang2024e5vuniversalembeddingsmultimodal). An adaptation of the E5 embedding approach to multimodal models: it trains only on text pairs yet bridges the modality gap to handle image inputs, reducing cost while achieving universal embeddings.
ColPali (faysse2025colpaliefficientdocumentretrieval). A vision-based document retrieval model that processes document pages as images (no OCR) and produces multi-vector embeddings via a late-interaction mechanism over PaliGemma (beyer2024paligemmaversatile3bvlm), enabling efficient and accurate retrieval.
ColQwen2.5 (faysse2025colpaliefficientdocumentretrieval). An extension of ColPali (faysse2025colpaliefficientdocumentretrieval) using Qwen2-VL (wang2024qwen2vlenhancingvisionlanguagemodels) as the backbone, carrying forward the late interaction retrieval paradigm over page image embeddings, capturing layout and textual context without OCR.
Jina-v4 (günther2025jinaembeddingsv4universalembeddingsmultimodal). A multimodal embedding model combining visual and textual inputs with support for multi-vector (late interaction) embeddings, using adapters over a unified backbone to excel on visually rich document retrieval.
NemoRetriever (xu2025llamanemoretrievercolembedtopperforming). An LI retriever that combines vision-language embeddings with a ColEmbed design, enabling high performance on visual document retrieval with structured patch matching and efficient similarity.
Jina CLIP (koukounas_jina_2024). A smaller scale vision-language model using CLIP embeddings, applied to document retrieval tasks; although not LI, it offers a lightweight multimodal baseline.
BGE Visualized M3 (zhou2024vistavisualizedtextembedding). A vision-enhanced version of BGE M3 (chen_bge_2024) that supports visual inputs and extends embedding models into multimodal domains.
SigLIP2-L-512/16 (tschannen2025siglip2multilingualvisionlanguage). A multilingual vision-language bi-encoder model, which combines image and text modalities to yield unified embeddings across languages. This configuration handles images of 512x512 pixels and create subpatches of 16x16 pixels.
ColFlor (masrycolflor). A lightweight OCR-free visual document retriever with only 174M parameters built over Florence-2 and DaViT, delivering strong performance near ColPali with much lower computational cost and much faster encoding.
@@ Appendix C Additional Ablations
C.1 Performance Against Off-the-Shelf Dual Encoder
We study whether using off-the-shelf performances of the standalone vision tower are not outweighing the burden of adding language parameters and re-training through language modeling, as proposed in our work. Figure 7 shows the results of the various models on the tasks described in Section 2. Similarly to Section 3.1, we observe that the early fusion model trained with LM objective significantly outperform the standalone vision tower on document retrieval tasks (+10.9 nDCG@5). It even surpass the larger dual encoder (+4.8 nDCG@5) on these latest tasks. We note that the standalone vision tower largely outperform the early fusion models on the other natural images tasks, supporting for the use of the SigLIP model for these tasks as found in various general benchmarks (xiao2025miebmassiveimageembedding).
C.2 Scaling Dynamics of Attention Masks
@@
C.3 Bridging the Gap with Longer Contrastive Training
@@
C.3.1 Optimal Text-To-Image Ratio for Document Retrieval
@@
C.4 Late Interaction for Non-Documental Retrieval
@@
We want to study if the multi-vector gains transfer to non-documental retrieval. To do so, we contrastively post-train our base model on our generalist post-training mix presented in Table 7. The late interaction generalist exhibits superior performance in retrieval setting, improving its single-vector performance by +20.2% (11.5 points), matching the performance of substantially larger VLM-based retrievers like E5-V (8.3B parameters, 67.5 points) and surpassing dual encoders like SigLIP (882M parameters, 56.7 points). This matches the capabilities observed in Section 3.1 for documental settings for models with native bidirectional attention, extending it to natural image tasks. This result extends the prevailing understanding from the document retrieval community, where the superiority of late-interaction is well-documented (khattab_colbert_2020, GTE-ModernColBERT, faysse2025colpaliefficientdocumentretrieval). While this performance gap is widely accepted for document retrieval, its applicability to caption matching tasks has not really been addressed. Our findings provide strong evidence that the fine-grained matching capabilities of late-interaction models are a key driver of performance in this domain too.
C.4.1 Model Merging
Our contrastive learning stage provides direct performance trade-offs on different tasks. Following recent trends, we evaluate how model merging techniques allow to mitigate performance degradation on specific tasks, while maintaining the performance enabled by the contrastive training (sung2023empiricalstudymultimodalmodelmerging; dziadzio2024mergemultimodalmodelstime; li2024improvinggeneraltextembeddingmerging; zhang2025qwen3embeddingadvancingtext). We merge our ablation model after modality alignment with the checkpoint after the full contrastive learning with two methods: SLERP (ilharco2022patchingopenvocabularymodelsinterpolating) and average merging (slerp_original). For SLERP, we compare three values for the $\lambda$ coefficient ($0.25$, $0.5$, $0.75$). Figure 11 displays the the trends with the best method (SLERP, $\lambda=0.75$). As we can see, the merged model mitigates the performance drop in Image/Caption Retrieval tasks, while maintaining significant gains on Image Classification tasks. However, merging strongly degrades performance on Document Retrieval, showing that benefits of merging embedding models are task-dependent.
@@ C.4.1 Model Merging
C.4.2 Curriculum For Document Retriever Contrastive Post-Training
@@
C.5 Text-Only Retrieval
@@
C.6 Model Latency
C.6.1 Image Resolution Tradeoffs
Figure 12 presents the pixel shuffling trade-off. Processing larger images creates more visual tokens, leading to very long sequences (around $17^{\prime}500$ tokens for a 2048x2048 px image with no pixel shuffling). Pixel shuffling allow to compress these sequence by concatenating the embeddings of spatially close patches. This diminishes the number of tokens for longer visual token embeddings. Table 11 presents the latency to process one image of various resolutions on one L4 GPU and CPU.
Figure 12: Image processing parameters impact on visual tokens. Here we assume a square image for simplicity. Scaling the image size introduces naturally more tokens, but having a large enough pixel shuffling ratio ($r\geq 4$) allows to counterbalance by concatenating spatially close patch representations.
Num. Visual Tokens CPU Latency (ms) GPU Latency (ms)
512512px $128$ $287.2_{(\pm 7.8)}$ $43.6_{(\pm 1.4)}$
10241024px $320$ $1015.8_{(\pm 58.1)}$ $150.3_{(\pm 2.5)}$
20482048px $1088$ $2572.0_{(\pm 63.9)}$ $363.4_{(\pm 4.6)}$
Table 11: ModernVBERT image processing latency. Computing the average time to process a single image on GPU and CPU. The average is computed on 100 images. The values represent the mean latency in milliseconds, with the standard deviation included in parenthesis.
C.6.2 Online Query Encoding Latency
We evaluate the query embedding speed of our model on GPU. We use a single Nvidia H100 with 80GB of VRAM. As for Section 4.2, latencies are computed in batch size 1 to simulate online situations, and are averaged over all NanoBEIR queries. Only the text parameters are loaded and run, to minimize memory usage. Parameters are cast to bfloat16 and Flash Attention 2 is used. The resulting speeds are often much faster than those obtained by running inference through each model’s reference implementation. Results are shown in Table 12). Interestingly in this setup where memory is not a bottleneck, model depth seems to be a large performance driver, sometimes more the parameter count. We finally evaluate batched GPU throughput. We use batches of size 512 by default and iteratively half it when memory is insufficient. We observe that ModernVBERT based models are extremely fast and can process 5000 queries per second. In the table, the reported figures correspond to the inverted throughput (latency per batch divided by the number of queries per batch). These speed and throughput gains are made possible due to a combination of size, and efficient hardware-informed design as well as the support of flash attention and sequence packing other models of the size often lack (warner2024smarterbetterfasterlonger).
Late Interaction Model Size (B)
CPU Latency (ms)
GPU Latency (ms)
GPU Batching (ms)
$\geq$ 1B Parameters
MoCa-3B 3.75 $158_{(\pm 147)}$ $26_{(\pm 3)}$ $4.54$
VLM2Vec 4.15 $211_{(\pm 253)}$ $21_{(\pm 3)}$ $2.82$
GME-Qwen2-7B 8.29 $412_{(\pm 411)}$ $25_{(\pm 1)}$ $9.07$
E5-V 8.36 $434_{(\pm 379)}$ $22_{(\pm 2)}$ $9.55$
ColPali 2.92 $175_{(\pm 113)}$ $14_{(\pm 1)}$ $3.07$
ColQwen2.5 3.75 $158_{(\pm 147)}$ $_{(\pm 2)} 2626$
Jina-v4 3.75 $158_{(\pm 147)}$ $26_{(\pm 2)}$ $4.54$
NemoRetriever-3B 4.40 $155_{(\pm 118)}$ $20_{(\pm 2)}$ $4.59$
$\leq$ 1B Parameters
Jina CLIP .22 $14_{(\pm 7)}$ $6_{(\pm 2)}$ .69.69
BGE Visualized M3 .87 $38_{(\pm 42)}$ $10_{(\pm 2)}$ .77.77
SigLIP2-L-512/16 .88 $25_{(\pm 8)}$ $6_{(\pm 1)}$ .10.10
ColFlor .17 $17_{(\pm 9)}$ $8_{(\pm 2)}$ .31.31
BiModernVBERT (ours) .25 $20_{(\pm 11)}$ $14_{(\pm 2)}$ .20.20
ColModernVBERT (ours) .25 $20_{(\pm 11)}$ $14_{(\pm 2)}$ .20.20
Table 12: Text query encoding latency. The latency is computed both on high-end CPUs (1TB RAM, 128 cores) and GPU (Nvidia H100, 80GB) (mean ± std). Since only 649 queries are used, standard deviations are not reported in GPU batching mode (batches of 512 queries by default), for which we report the inverse throughput (average latency per batch divided by the batch size).